OpenStudy (anonymous):

Evaluate the expression when x = –24. A. –6 B. –2 C. 2 D. 6

Mehek (mehek14):

expression?

OpenStudy (kweenlexis):

whats the expression?

Mehek (mehek14):

substitute x = 24 \(\dfrac{24-12}{6}\\24-12=12\\ \dfrac{12}{6}=2\)

OpenStudy (kweenlexis):

yeah what @Mehek14 :)

OpenStudy (kweenlexis):

said*

OpenStudy (anonymous):

C?

Mehek (mehek14):

yea

OpenStudy (anonymous):

can u help me with 1 more

Mehek (mehek14):

sure :]

OpenStudy (anonymous):

What is the mean of the set of numbers? {–11, 22, –2, –3} A. –6.5 B. –1.5 C. 1.5 D. 6.5

Mehek (mehek14):

add all of the numbers

Mehek (mehek14):

then divide by 4

OpenStudy (kweenlexis):

are you in 6th grade or 5th?

OpenStudy (anonymous):

B or C?

OpenStudy (anonymous):

i think C

Mehek (mehek14):

-11 + 22 = 11 11 - 2 = 9 9 - 3 = 6 6/4 = 1.5

Mehek (mehek14):

so yes C

OpenStudy (anonymous):

can u help with reading

OpenStudy (anonymous):

or writing

Mehek (mehek14):

I'm not good in writing :(

Mehek (mehek14):

you can tag people in that section

OpenStudy (anonymous):

like this

OpenStudy (muscrat123):

maybe i can help w the writong

OpenStudy (anonymous):

Is the underlined word a participle or verb? Are you cooking a special birthday dinner? A. participle B. verb

OpenStudy (muscrat123):

*writing

OpenStudy (muscrat123):

which word is underlined?

OpenStudy (kweenlexis):

which word is underlined?

OpenStudy (anonymous):

cooking

OpenStudy (muscrat123):

its a verb

OpenStudy (kweenlexis):

i think it would be B too.

OpenStudy (muscrat123):

b/c a verb is an action

Mehek (mehek14):

oh english I can help with not like writing paragraphs

OpenStudy (muscrat123):

lol

OpenStudy (anonymous):

Is the underlined word a participle or verb? There was standing water in the basement after the flood. A. verb B. participle standling is underlined

Mehek (mehek14):

and yes cooking = verb

Mehek (mehek14):

standing = verb

OpenStudy (kweenlexis):

which is the underlined word?

Mehek (mehek14):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @shotgun Is the underlined word a participle or verb? There was standing water in the basement after the flood. A. verb B. participle standling is underlined \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (muscrat123):

it said standing

Mehek (mehek14):

it says the word :P

OpenStudy (anonymous):

Is the underlined word a participle or verb? The troops marched across the field wearing (tattered) uniforms. A. verb B. participle

OpenStudy (kweenlexis):

oh i aint read that my bad

OpenStudy (muscrat123):

well its not a verb

Mehek (mehek14):

^

OpenStudy (muscrat123):

bc tattered is not an action

OpenStudy (muscrat123):

any more ?

OpenStudy (anonymous):

6 more

OpenStudy (anonymous):

Is the underlined word a participle or verb? The kids are polishing their bikes for the parade. A. verb B. participle

OpenStudy (candyme):

On that first problem you did, make sure you substituted -24. It looks like you guys put in a positive 24.

OpenStudy (anonymous):

polished

OpenStudy (muscrat123):

verb bc it is an action

OpenStudy (muscrat123):

and yea @candyme is right its -24 not 24

Mehek (mehek14):

so -2

OpenStudy (anonymous):

Is the underlined word a participle or verb? The kids are polishing their bikes for the parade. A. verb B. participle

OpenStudy (anonymous):

sorry lol

OpenStudy (candyme):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mehek14 substitute x = 24 \(\dfrac{24-12}{6}\\24-12=12\\ \dfrac{12}{6}=2\) \(\color{#0cbb34}{\text{End of Quote}}\) SO this should have looked like \[\frac{ -24-12 }{ 6 }\]\[\frac{ -36 }{ 6 }\] Giving you a -6, not 2